### v1

@@ Appendix C Circulation regime bistability
The fact that the circulation regime changes not when the mesh is stretched (comparing UniformParam to StretchParam) but when the convection scheme is disabled (comparing StretchParam to StretchExplicit), confirms our key finding that for the global climate mesh stretching is less important than changes in model parameterizations (see Sec. 3). Furthermore, our StretchExplicit experiment is qualitatively similar to the convection-permitting simulation of TRAPPIST-1e in Yang et al. (2023), in terms of the spatial distribution of temperature, wind, cloud cover, and precipitation. This implies that at least some of the differences between the coarse-grid GCMs and the convection-permitting model reported by Yang et al. are due to the circulation regime change and not due to resolving convection per se. And as Sergeev et al. (2022a) demonstrated, the circulation regime change can happen even in a coarse-grid GCM if the convection parameterization is switched off or modified.

### v2

@@ Appendix C Circulation regime bistability
The fact that the circulation regime changes not when the mesh is stretched (comparing UniformParam to StretchParam) but when the convection scheme is disabled (comparing StretchParam to StretchExplicit), confirms our key finding that for the global climate mesh stretching is less important than changes in model parameterizations (see Sec. 3). Furthermore, our StretchExplicit experiment is qualitatively similar to the convection-permitting simulation of TRAPPIST-1e in Yang et al. (2023), in terms of the spatial distribution of temperature, wind, cloud cover, and precipitation. This implies that at least some of the differences between the coarse-grid GCMs and the convection-permitting model reported by Yang et al. are due to the circulation regime change and not due to resolving convection per se. And as Sergeev et al. (2022a) demonstrated, the circulation regime change can happen even in a coarse-grid GCM if the convection parameterization is modified or switched off.
Despite the circulation regime change in these simulations, our model predicts a climate within the spread of THAI GCM results (Sergeev et al., 2022b). Taking the experiment with the uniform mesh (UniformParam) as an example, the global mean surface temperature is (Fig. 7h), which is lower than that predicted by ExoCAM and LMD-G but higher than that predicted by the UM for the Hab 1 case. Likewise, the maximum surface temperature () falls within the inter-model spread. The lowest surface temperature is , simulated in the regions of night-side stationary gyres. It is on the colder end of the spectrum of the THAI GCMs but still about higher than that in the UM simulation. Comparing the StretchExplicit case with the ROCKE-3D simulation in Sergeev et al. (2022b), we see the same trend in the surface temperature in both models: the minimum temperatures rise to about due to the shift of the stationary waves, while the maximum temperatures fall to . Thus, in the THAI Hab 1 case, LFRic-Atmosphere tends to produce a colder climate compared to other GCMs, though not as cold as that in the UM.